A A mechanic is paid $20 per hour for regular time, the first 40 hours, and time-and-a-half for overtime, with maximum of 60 hours week. Let h be the number of hours worked and W be the weekly wage for the mechanic.
a) Write a piecewise-defined function for the hourly wage h and a piecewise-defined function for the weekly wage of the mechanic.

Answer by Theo(13342)

h = number of hours worked.
w = weekly wage for the mechanic.

the piece wise defined function would probably be defined as:

w = 20 * h for h <= 40 + 30 * (h - 40) for 40 < h <= 60.

another way to do it would be:

w = 20 * h for h <= 60 + 10 * h for 40 < h <= 60.

for example:

person worked 40 hours.

pay would be 20 * h in either formula.

person worked 60 hours.

pay would be 20 * 40 for the first 40 hours and 30 * 20 for the hours worked over 40 in the first formula.
total pay would be 40 * 20 + 20 * 30 = 1400.

pay would be 20 * 60 for the 60 hours plus 10 * 20 for the hours worked over 40 in the second formula.
total pay would be 60 * 20 + 20 * 10 = 1400.

either way, the person gets payed 20 dollars an hour for the first 40 hours plus 30 dollars an hour for any hours greater than 40 but less than or equal to 60.